OpenStudy (anonymous):

A line passes through the point (7,-4_ and has a slope of 2 . Write an equation for this line.

OpenStudy (anonymous):

Same logic i used on the other question. y = mx + b the slope (m) is given and its = 2 you know the point (7,-4) so now all you must do is to find b good luck

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

well.. To find b you need to add all you have to the equation: y = mx+b -4 = 2*7 + b b = -4 - 14

OpenStudy (anonymous):

lol \[y-y_1=m(x-x_1)\] \[y+4=2(x-7)\] solve for \(y\) in two steps \[y+4=2x-14\] \[y=2x-18\]

OpenStudy (anonymous):

@satellite73 didn't knew that technique. Awesome stuff!!!